OpenStudy (anonymous):

An urn contains 9 white balls and 10 red balls. Three balls are selected. In how many ways can the 3 balls be drawn from the total of 19 balls: a) If 2 balls are white and 1 is red? b) If all 3 balls are white? c) If all 3 balls are red?

OpenStudy (anonymous):

same as the last one

OpenStudy (anonymous):

two white, one red \[\binom{9}{2}\times \binom{10}{1}\] which is really just \[\frac{9\times 8}{2}\times 10\]

OpenStudy (anonymous):

3 white \[\binom{9}{3}\] 3 red \[\binom{10}{3}\]

OpenStudy (anonymous):

so do you always just subtract 1 for the like 9x8 thing (like where did you get 8?)

OpenStudy (anonymous):

lets compute 9 choose 2 by the counting principle

OpenStudy (anonymous):

9 choices for the first ball then there are 8 left, so 8 choices for the second ball by the counting principle, the number of ways to do that is \(9\times 8\) then you have to divide by 2 because the above method counts \((a,b)\) differently from \((b,a)\)

OpenStudy (anonymous):

similarly \[\binom{9}{3}=\frac{9\times 8\times 7}{3\times 2}\]

OpenStudy (anonymous):

ok! so is it a) 360 b) 84 c) 120

OpenStudy (anonymous):

\[\binom{12}{4}=\frac{12\times 11\times 10\times 9}{4\times 3\times 2}\] sometimes written as \[\frac{12\times 11\times 10\times 9}{4!}\]